OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y greater than or equal to −4x − 2? Is the answer (-1,0) ?

OpenStudy (amistre64):

try it out 0 >= -4(-1)-2 0 >= 4-2 0 >= 2 is 0 bigger than or equal to 2?

OpenStudy (anonymous):

im not sure , i do know that 0 is greater than -2

OpenStudy (anonymous):

0 is less than 2 thou @amistre64 .. so id be incorrect.. the answer would be (0,0) ?

OpenStudy (amistre64):

try it out 0 >= -4(0)-2 0 >= 0-2 0 >= -2 that is one solution yes.

OpenStudy (anonymous):

it was multiple choice (−1, 0) (0, 0) (4, 4) (1, 3)

OpenStudy (anonymous):

i was stuck between -1,0 and 0,0

OpenStudy (amistre64):

:) then id definanlty go with 0,0

OpenStudy (anonymous):

thanks